The wavelength of a distant star is 5700A0 and the spectral light has a shift of 1.9A0 towards red end then the velocity of star relative to the earth will be?

Grade:12th pass

1 Answers

Pankaj
4 years ago
∆lambda = (V/C)×lambda
V = velocity of planet
C = velocity of light
∆lambda = change in wavelength
Lambda = original wabelwavel
Putting values, 
1.9= (V/3×10^8)×5700
V = 1×10^5 m/sec